OpenStudy (anonymous):

condense ln 4+5 ln x-ln y

OpenStudy (whpalmer4):

\[\log a+ \log b = \log (a*b)\]\[\log a - \log b = \log (\frac{a}{b})\]

OpenStudy (whpalmer4):

Also \[a\log b = \log (b^a)\]